b. Determine whether each of these functions is a bijection from Z to Z.

f (n) = n2 + 1



1
Expert's answer
2021-05-07T10:22:54-0400

Let us determine whether the function "f:\mathbb Z\to\mathbb Z," "f (n) = n^2 + 1" is a bijection. Since for "-1\ne 1" we have that "f(-1)=2=f(1)", we conclude that the function "f" is not an injection. Consequently, the function "f" is not a bijection.